OpenStudy (anonymous):

Solve using elimination: -4y + 2x − 8 = 0 -3x + 6y = 12 Infinitely many solutions No solution (3, -1) (-1, 4)

OpenStudy (anonymous):

3(-4y + 2x − 8 = 0) 2(-3x + 6y = 12) -12y+6x-24=0 -6x+12y=24 cancel the x's and y's

OpenStudy (anonymous):

IS IT THE THIRD ONE